### v1

Abstract
In this paper, we investigate the prospects and challenges of sensor suites in achieving autonomous control for flying insect robots (FIRs) weighing less than a gram. FIRs, owing to their minuscule weight and size, offer unparalleled advantages in terms of material cost and scalability. However, their size introduces considerable control challenges, notably high-speed dynamics, restricted power, and limited payload capacity. While there have been notable advancements in developing lightweight sensors, often drawing inspiration from biological systems, no sub-gram aircraft has been able to attain sustained hover without relying on feedback from external sensing such as a motion capture system. The lightest vehicle capable of sustained hover—the first level of “sensor autonomy”—is the much larger 28 g Crazyflie. Previous work reported a reduction in size of that vehicle’s avionics suite to 187 mg and 21 mW. Here, we report a further reduction in mass and power to only 78.4 mg and 15 mW. We replaced the laser rangefinder with a lighter and more efficient pressure sensor, and built a smaller optic flow sensor around a global-shutter imaging chip. A Kalman Filter (KF) fuses these measurements to estimate the state variables that are needed to control hover: pitch angle, translational velocity, and altitude. Our system achieved performance comparable to that of the Crazyflie’s estimator while in flight, with root mean squared errors of 1.573 degrees, 0.186 m/s, and 0.139 m, respectively, relative to motion capture.
I INTRODUCTION
Insect-scale robots have the potential to be deployed in the thousands or millions to perform “fast, cheap, and out of control” space missions, collective assembly tasks, or hazard detection, owing to their small size and low materials cost. Recent incarnations of sub-gram aircraft have demonstrated controlled flight [1, 2], as well as electromagnetically-mediated power delivery [3, 4, 5]. But sensor autonomy, that is, the ability to hover without external feedback, has not yet been demonstrated below 10 g, never mind 1 g. A central question is “What are the minimum sensor suite and computation resources needed for the task of flight control?” Of particular concern when designing such a control architecture is how the physics of scale affects sensing and control. The paper [6] serves as a step toward a missing but complementary analysis to previous work that has largely focused on scaling effects on actuation and mechanics in small robots [7, 8, 9, 10].
Insect-scale robots have extreme size, speed, weight, and power (SSWaP) constraints. “Sensor autonomy” including stable hovering, represents first level of autonomy in the hierarchy proposed in [11], and is the foundation upon which higher level capability is built. We impose a requirement that sensing and computation to take no more than 10% of the power required to stay in the air; any more than this and flight time is severely impacted. This factor holds for important examples of autonomous drones, such as the 1.5 kg system in [12] and the 30 g system in [13]. This suggests a power budget of $\sim$20 mW for a 150 mg Robofly (Figure 1) that consumes 100 mW to fly (about 200 mW after losses from a 50% efficient boost converter are factored in).
Figure 1: The TinySense sensor suite shown next to a U. Washington Robofly and a standard pencil for scale. The width of the sensor suite board is approximately 12 mm and its mass is 78.4 mg.
The sensing system is very crucial for robots. Large drones can have more payloads to hold sensors onboard such as inertial measurement units (IMUs) and the global positioning system (GPS). However, GPS does not work well in indoor environments [14] and small robots are not able to hold many sensors due to constraints in size, speed, weight, and power consumption (SSWaP). There are several work done on the sensor suite for insect-scale robots  [15],  [6],  [16]. The most recent sensor suite is 187 mg and consumes 21 mW [15].
In this manuscript, we introduce a new avionics system for autonomous hovering flight that is even better tailored in mass and energy consumption for an insect-scale robot (Fig. 3) than previous work [15]. To do so, we made two key refinements to lower power and mass. First, we replaced the power-hungry laser rangefinder with a pressure sensor. In recent years these have attained sufficient precision to be used for the purpose of altitude estimation and control even for insect-scale robots. Second, we replaced the optic flow sensor with a global shutter camera and compute optic flow on a microcontroller, resulting in a reduction in weight from 97 mg to 44 mg. As in [15], this new sensor suite is able to estimate the crucial variables that are needed for controlled flight: pitch angle, translational velocity, and altitude.
Our contributions, in order of importance, are:
  1. 1.
    A significant reduction in weight of the sensor suite to 78.4 mg which is less than half of the previous lightest sensor suite [15].
  2. 2.
    We showed that we were able to make accurate estimates of the state variables pitch angle, translational velocity, and altitude using data collected wirelessly from this sensor suite while it was mounted on a flying helicopter.
  3. 3.
    Implemented an optic flow algorithm on a microcontroller.
Subsequent sections elucidate the dynamics, sensor configurations and measurement models, and corroborate our findings with data acquired wirelessly from our sensor suite, which is then analyzed on a desktop computer.
II METHODOLOGY
II-A Dynamics
The basic dynamics of most small hovering devices are unstable in pitch angle and position. This was true for the first robotic flies [17] and EHD thruster actuated robots [18], as well as larger aircraft such as the tailless Delfly Nimble [19]. Techniques to stabilize them passively through modification of wing configuration [20] or using air dampers [21] are either unproven or result in undesirable steady-state oscillations, respectively. To stabilize these dynamics, we will use a linearized dynamical model from which we will construct an observer of these unstable states so that they can be controlled.
Figure 2: Forces and state variables of a small hovering Flying Insect Robot (FIR).
Figure 2 depicts forces acting on a flapping-wing robot such as a gnat robot or the Robofly, but the proposed model can apply to almost any small flying aircraft, including EHD-based thrusters [22].
Near hover, the vehicle’s dynamics are well approximated by a linear state-space model [23, 24, 21] of the form $\dot{\bm{q}}=A\bm{q}+B\bm{u}+G\bm{w}$, where $\bm{w}$ is the process noise, assumed to be roughly zero-mean Gaussian white noise with known covariance. Process noise $\bm{w}$ is included because it is a necessary consideration for the design of the estimator.
To simplify the analysis, we restrict our attention to motion in the $x$-$z$ plane; the dynamics in the $y$-$z$ plane are very similar. We choose a minimal state vector $\bm{q}=[\theta,v_{x},z]^{T}$, where $\theta$ is the pitch angle, $v_{x}$ is the vehicle’s translational velocity in the world $x$-direction, and $z$ is its vertical position or altitude (Fig. 2). When the nonlinear Euler-Lagrange equation dynamics are linearized by taking the Jacobian at $\theta=0$, thrust force $f_{t}=mg$, and $z=z_{d}$, where $z_{d}$ is the desired altitude, the dynamics matrices are given by:
$\footnotesize A=\left[\begin{matrix}0&0&0\\ g&-\frac{b}{m}&0\\ 0&0&0\end{matrix}\right],\quad B=\left[\begin{matrix}1\\ 0\\ 0\end{matrix}\right]$ (1)
, where $m$ is the vehicle’s mass, and $b$ is the vehicles translational drag proportionality constant [16]. Note that because we are concerned in this paper about only estimation, we ignore inputs from the motor, and instead use the measurements of the gyroscope $\bm{u}=[\omega_{m}]$ as the only input. This is discussed further in Section V.
III Sensor Suite and Measurements
Next we consider the sensor suite and state estimator. Our interest is in sensors that are either available commercially off-the-shelf, nearly so, or can be fabricated using the same tools that have been successful building insect robots such as the Robofly [25] and Robobee [26, 1]. A further constraint is that they require extremely low power for both the sensor itself and subsequent transduction and processing. This largely eliminates sensors that emit power, such as radar, sonar, depth cameras that emit structured light, and scanning laser rangefinders [27], but permits passive sensing such as vision. The global positioning system is largely denied indoors and does not provide enough resolution anyway (1–10 m), and, further, requires significant power for signal processing. A lightweight and low-power sensing modality that can provide the necessary information to hover is “optic flow,” which is defined as the velocity of the image (in rad/s) across the camera imager [28]. In tandem with other sensors, optic flow can be used to estimate the distance to obstacles as the imager translates through space [29]. Methods for estimating optic flow include calculating luminance gradients locally [30] (Lucas-Kanade), or image-wide [31], as well as iterative search for matching blocks of pixels [32]. Distances can also be estimated with a second camera and stereo vision using line matching [33], but this requires a second camera and a long “baseline” between cameras for accuracy, which would add too much weight.
Figure 3: (a) Sensor suite (PAG7920LT optical sensor, ICM42688-P IMU, BMP388 pressure sensor) next to penny.
Given these constraints, we propose the following set of onboard sensors: An optical sensor, a pressure sensor, and a gyroscope. We designed a customized microcircuit with the three sensors. (Fig. 3)
III-A The camera and optic flow
The system employs an optical sensor, the PAG7920LT (PixArt Imaging, Inc., Taiwan), to capture $160\times 120$ QQVGA images at 100 Hz over a 32 MHz SPI connection, with which the microcontroller computes optic flow via the Lucas-Kanade Method. The reason we select this camera is that, to our knowledge, it is the only global shutter camera that provides an SPI interface to transmit the image data. Other global shutter cameras require a special interface called MIPI (mobile industry processor interface), which is not available for any microcontroller. The majority of cameras employ a “rolling shutter” in which pixel rows are captured at sequential periods of time. A global shutter camera, by contrast, captures that all pixels at the same time, which ensures an accurate optic flow measurement.
To minimize the data transmitted wirelessly over Bluetooth from our TinySense sensor suite, it is necessary to perform all optic flow calculations onboard. We wrote a Lucas-Kanade algorithm using the Arduino development environment in C++ to run on NRF52-based Adafruit Express board, chosen for its low size and weight.
The Lucas-Kanade algorithm is performed by calculating the following quantities:
$A=\left[\begin{matrix}I_{x}(q_{1})&I_{y}(q_{1})\\ I_{x}(q_{2})&I_{y}(q_{2})\\ ...&...\\ I_{x}(q_{n})&I_{y}(q_{n})\end{matrix}\right]\bm{v}=\left[\begin{matrix}V_{x}\\ V_{y}\end{matrix}\right]b=\left[\begin{matrix}-I_{t}(q_{1})\\ -I_{t}(q_{2})\\ ...\\ -I_{t}(q_{n})\end{matrix}\right]$
, where $q_{1},q_{2},\dots,q_{n}$ are the brightness readings of each of the $n$ pixels in the image, $\bm{v}$ is the image flow in pixels per frame, and $I_{x}(q_{i}),I_{y}(q_{i}),I_{t}(q_{i})$ are the image’s horizontal, vertical, and temporal (time) partial derivatives respectively, at pixel $q_{i}$. Then, the optic flow is estimated (in pixels/frame) by taking the pseudoinverse $\bm{v}=(A^{T}A)^{-1}A^{T}b$.
To minimize computation for the limited microcontroller resources, we reduced the image resolution by a factor of 4 (producing a $40\times 30$ resolution) by skipping pixels.
The two most expensive operations involve computing $A^{T}A$ and $A^{T}b$. A $40\times 30$ image produces an $A$ matrix with $(40-2)*(30-2)=1064$ rows and 2 columns and a $b$ matrix with 1064 rows and 1 column. We estimate multiplying two matrices of size $k\times l$ and $l\times m$ will require $kml$ multiplications. Thus, computing $A^{T}A$ and $A^{T}b$ require 4256 and 2128 multiplications, respectively. We neglect the additional operations required to compute their product (2 multiplications) and the $2\times 2$ matrix inverse (6 multiplications and one divide), giving a total operations count of 6384.
We additionally implemented a form of sparse optic flow in which flow vectors were computed for twelve $10\times 10$ patches, which are averaged to produce a final estimate. A $10\times 10$ image produces an $A$ matrix with $(10-2)*(10-2)=64$ rows and 2 columns and a $b$ matrix with 64 rows and 1 column. Thus, computing $A^{T}A$ and $A^{T}b$ require approximately 256 and 128 multiplications, respectively. With twelve matrices of size $10\times 10$, the total number of multiplication is $12(256+128)=4608$. This is a reduction of approximately $39\%$. In the future, we will evaluate these patches individually to remove outliers, removing patches with low eigenvalues (low contrast) or significantly different eigenvalues (aperture problem).
To convert the image flow in pixels/frame to angular velocity in radians/s, we rotated the TinySense by hand and minimized the error between the optic flow measurement multiplied by a constant $\alpha$ and the gyroscope reading. This constant, which depends on the pixel density and field of view of the camera, is given by $\alpha=-0.027$ rad/pixel.
III-B Pressure sensor
Altitude measurements are conducted by a Bosch BMP388 pressure sensor. This sensor estimates altitude as a proportion of the difference between the surrounding air pressure and sea level. We set it for 2$\times$ pressure oversampling and an IIR filter coefficient of 3 to maximize resolution. To compensate for variations in local pressure causing inaccurate altitude measurements, the sensor was assigned a bias equal to the average of the first 25 measurements where the sensor system was grounded; this bias was subtracted from every reading.
III-C The gyroscope
For the onboard inertial measurement unit (IMU), we chose the TDK ICM42688-P because it is a recent model that is available in the smallest commercially-available package size of 2.5 $\times$ 3 mm. It includes both a 3-axis gyroscope and a 3-axis accelerometer. In this work, we used only the gyroscope, which measures angular velocity through the detection of Coriolis forces within an electromechanical resonator. We configured the sensor to have a ±$\pm 2000$ degrees/s full-scale (FS) range and enabled the sensor-provided low-noise mode to improve the sensitivity and minimize noise. During initial experiments, we noticed that the gyroscope’s signal was noisier than expected and differed significantly from the Crazyflie’s gyroscope, even after applying a digital low-pass filter. In handheld, non-flying experiments, however, the gyroscope proved accurate. We hypothesized that vibrations from the spinning propellers, which were at a higher frequency than the 100 Hz sampling rate, were introducing aliasing. We enabled the sensor’s anti-aliasing low-pass filter with a cutoff frequency of 42 Hz, which is well below the Nyquist frequency associated with the 100 Hz sampling rate. We found that this setting significantly reduced the gyroscope’s noise and error.
TABLE I: Comparison between TinySense and prior work
component size mass power
(mm) (mg) (mW)
prior work [15] IMU 2.5$\times$3$\times$0.91 14 3
rangefinder 4.9$\times$2.5$\times$1.56 16 5
optical flow 5$\times$5$\times$3.08 97 12
board/discretes 60
total 187 21
TinySense IMU 2.5$\times$3$\times$0.91 14 3
pressure 2.0$\times$2.0$\times$0.75 14 2.6
camera chip 2.2$\times$2.7$\times$6.9 24 10
lens 3.8$\times$3.8$\times$1.95 20
board/discretes 26.4
total 78.4 15.6
IV State estimation
To implement a Kalman Filter, we must first ascertain whether our system is observable at equilibrium. To show this, we linearize the observation model for the task of hovering. Note that the gyroscope provides a low-noise measurement of angular velocity $\bm{\omega}$. In cases like this, where the sensor provides more precise information than can be derived from knowing the system’s actuated inputs (e.g. the torque applied by the aircraft’s rotors or wings), it is common practice to re-formulate the Kalman Filter to use the sensor’s readings as its "input" $\bm{u}$, rather than the actual system input. Doing this has the additional advantage that for each such measurement, one less state variable is required, reducing computation requirements and complexity. In this case, it means that the angular velocity $\omega$ need not appear among the state variables. The following is our observation model for the suite of sensors, which is:
$\bm{y}=\left[\begin{matrix}\Omega_{m}+n_{o}\\ z_{m}+n_{p}\\ \end{matrix}\right],$ (2)
where $\Omega_{m}$ is the optic flow measurements from optic flow sensor, $z_{m}$ is the measurement from the pressure sensor, and $n_{o}$ and $n_{p}$ are the noise terms for the two sensors. We assume noise terms are zero-mean Gaussian white noise, a standard assumption in sensor noise modeling to simplify the analysis. The measurement model for the optic flow in the $x$ direction, as measured by the optic flow camera is a nonlinear function of state [6]:
$\Omega_{m}=\omega_{m}-\frac{v_{x}}{z},$ (3)
where $\omega_{m}$ is the angular velocity measurement from gyroscope. We can see the optic flow measurement depends on $z$. We can take the Jacobian linearization at a desired altitude $z_{d}$:
$\begin{split}\Omega_{m}(z)=\Omega_{m}(z_{d})+O(z)=\omega_{m}-\frac{v_{x}}{z_{d}}+O(z).\end{split}$ (4)
With this linearization, our measurement model can be cast into the traditional Kalman Filter model, $\bm{y}=C\bm{q}+D\bm{u}+\bm{n}$, where
$C=\left[\begin{matrix}0&-\frac{1}{z_{d}}&0\\ 0&0&1\end{matrix}\right],\quad D=\left[\begin{matrix}1\\ 0\end{matrix}\right].$ (5)
The observability matrix [$C$; $CA$] is full rank, meaning that a Kalman Filter will be able to estimate the system’s state using the measurements given in Eq. (5).
The Kalman Filter observer estimates $\hat{\bm{q}}$ by numerically integrating the system
$\dot{\hat{\bm{q}}}=A\hat{\bm{q}}+B\bm{u}+K(\bm{y}-C\hat{\bm{q}}-D\bm{u}),$ (6)
where $K$ is the Kalman gain.
We used a gradient descent algorithm to minimize the RMSE between the TinySense state estimates and the mocap ground-truth to tune the process noise covariance matrix $Qn$ and sensor noise covariance matrix $Rn$:
$Qn=diag(0.14^{2},0.093^{2},0.08^{2}),\quad Rn=diag(1.06^{2},0.2^{2}).$
For the process noise (disturbance) matrix $G$, we assumed that the noise enters the system as white noise disturbance adding to the derivatives of the three states $\theta$, $v_{x}$, and $z$. Based on these, the Kalman gain $K$ was computed using the lqe command in python-control [34].
$G=\left[\begin{matrix}0.72&0&0\\ 0&0.10&0\\ 0&0&20.00\end{matrix}\right],\quad K=\left[\begin{matrix}0.00204&0\\ 0.48452&0\\ 0&1.81818\end{matrix}\right].$
We implemented the Kalman Filter offline in Python by iteratively computing the state estimates $\hat{\bm{q}}_{i+1}$ according to $\hat{\bm{q}}_{i+1}=\hat{\bm{q}}_{i}+\dot{\hat{\bm{q}}}_{i}\cdot dt$.
Figure 4: Sensor measurements from the onboard gyroscope, optic flow sensor, and pressure sensor of TinySense are compared with those from the Crazyflie for three different flight experiments (a), (b), and (c). In the altitude plots, altitude is measured by the Crazyflie’s laser rangefinder, while the TinySense uses a pressure sensor; mocap data is also included for comparison.
V Experiments
V-A Robot platform
We performed experiments using Crazyflie 2.0, a palm-sized quadrotor by Bitcraze (Sweden), in wind-free conditions in an indoor environment. In addition to its built-in gyroscope, the Crazyflie was equipped with the Flow deck v2 consisting of a downward-facing optic flow sensor and a laser rangefinder; it uses an Extended Kalman Filter to estimate its state in flight [35].
V-B Motion capture system
To provide a ground-truth comparison for the sensors, we used a twelve-camera OptiTrack Flex 13 motion capture arena. The Crazyflie drone was fitted with four retro-reflective motion capture markers and its position and orientation were recorded at 120 Hz. Velocity was estimated by taking the numerical derivative of the position data in post-processing.
V-C Experiment
We connected our sensor suite to an NRF52840 Express (Adafruit), a board which supports data transmission through UART and 2.4GHz Bluetooth Low Energy compatibility. We mounted the board as close to the geometrical center of the Crazyflie as possible to maintain a reasonable center of mass and minimize vibration interference to the IMU from the propellers. Power was supplied from the Crazyflie’s 250mAh Lithium Polymer (LiPo) battery.
Each trial consisted of commanding the robot to rise to an altitude of 1 m, translate forward 1 m at 1 m/s, and then descend back to the ground. We ran this experiment three times. We collected measurements from the TinySense’s sensors on a laptop as they were transmitted over Bluetooth at 100 Hz as binary data. Simultaneously, a second computer captured motion capture data at 120 Hz and used a Python-based script to capture sensor measurements and state estimates transmitted from the Crazyflie over Bluetooth at its maximum rate of 30 Hz. Each receiving device independently recorded a UTC timestamp, which was used to synchronize the data. The sensor measurements from the TinySense compared with the Crazyflie, and the motion capture (mocap) system’s measurement for altitude are shown in Fig. 4.
Figure 5: State estimates from TinySense after Kalman filtering, Crazyflie estimates, and mocap estimates are compared. We conducted three flight experiments (a), (b), and (c), with the Kalman filter estimates starting at time zero. Note that we ignore the pressure sensor measurements in the first 1.8s because of the pressure change near the surface.
TABLE II: Estimation of Optic Flow Algorithm Power Use
Update Frequency (Hz) Occurences/ Update Step Name Single-cycle Operations Int Division Operations Floating-point Division Operations Total Cycles/Update Total Cycles/Second Power Use
100 1 Skip frame 8493 0 8 249781 24.9781 Mhz $\sim$5.4552mW
12 LK OF 18434 128 2
1 Copy data 1200 0 0
V-D Computational Load
We estimated the power consumption of the optic flow algorithm from the number of cycles in a single run. We focus on only estimating three types of operations conducted on the ARM Cortex-M4 based microcontroller: single-cycle operations (addition, subtraction, multiplication, and memory read/write); integer divisions, which require 12 cycles at most; and floating-point divisions, which require 14 cycles. These results are summarized in Table II. Given that in our current system, all sensors update at a rate of 100 Hz, we gather that the total number of cycles per second is $\sim$24.98 Mhz. Operating on the nRF52840 chip with an efficiency of 52 $\mu$A/Mhz and a battery-powered voltage of 4.2V, we estimate the power use of the optic flow algorithm to be $\sim$5.45mW.
V-E Results and Discussions
Fig. 5 compares our sensor suite’s state estimate to that of the Crazyflie’s built-in estimator, and states derived from the mocap system, considered to be ground truth. For the state estimator, time $t=0$ corresponds to when the gyroscope reading becomes significantly different from zero, indicating liftoff. It can be observed in Fig. 4 that a pressure increase occurs when the propellers first turn on, which can be observed as a drop in measured altitude. To mitigate unwanted effects of this in the Kalman Filter, we ignored (set to zero) altitude measurements $z_{m}$ during the first 1.8s after liftoff. Nevertheless, once it is in the air, the altitude measurement matches quite closely with the laser rangefinder on the Crazyflie, though with slightly more noise.
Our system’s estimate compares well with those of the Crazyflie’s, despite receiving gyroscope readings at a much lower rate (100 Hz) compared to 1000 Hz on the Crazyflie[35]. The RMSE is approximately equal for pitch angle (Table III). We observe that there are occasionally large errors in translational velocity $v_{x}$. These appear to occur after sudden changes in attitude that occur around $t=5$ s and $t=6$ s, during which the angular velocity momentarily exceeds 3 rad/s (Fig. 4). We note that at these moments, our TinySense optic flow sensor’s readings are consistently lower than those of the Crazyflie. Lucas-Kanade cannot detect motion accurately if the displacement between frames is greater than a half pixel between image frames. The maximum allowable angular velocity is given by $\Omega=0.5\alpha F=(0.5)(0.027)(100)=1.4$ rad/s, where $\alpha$ is the angular increment between neighboring pixels (Section II-A) and $F$ is the frame rate in Hz. So the expected optic flow is faster than can be detected. This could be mitigated by a faster frame rate or a larger angular displacement between pixels.
TABLE III: Root Mean Squared Error (RMSE) of estimates relative to mocap system
velocity $v_{x}$ pitch angle $\theta$ altitude $z$
(m/s) (deg) (m)
Crazyflie 0.075 $\pm$ 0.009 1.619 $\pm$ 0.267 0.021 $\pm$ 0.001
TinySense 0.186 $\pm$ 0.015 1.573 $\pm$ 0.166 0.139 $\pm$ 0.026
VI Conclusion
In the present study, we introduce a new sensor suite with dramatically reduced mass, and associated optic flow and state estimation software that is compatible with the computational constraints of an onboard microcontroller small enough to fly aboard flying insect robots (FIRs) weighing less than a gram. Our proposed avionics suite diverges in two important ways compared to that of the 28 g Crazyflie helicopter, the lightest vehicle yet to perform controlled hovering, and previous work [15]. First, it replaces a power-hungry laser rangefinder with a much more efficient (and slightly less ideal) and lighter pressure sensor. Second, we replaced an off-the shelf optic flow sensor, which is available only in relatively heavy packages, with a custom camera with a global shutter and a custom-written Lucas-Kanade-based optic flow algorithm running onboard a 10 mg microcontroller small enough to fly onboard an FIR (estimated conpute power usage is about 5 mW). Our system uses a Kalman Filter to estimate pitch angle, translational velocity, and altitude. Remarkably, the cumulative weight of this sensor suite is only 78.4 mg, rather less than the 187 mg of previous work [15]. Incorporating a microcontroller to this configuration, the overall system remains well within the anticipated 252 mg payload limit of the 143 mg robotic platform reported in [18]. Our system demonstrates average RMSE values of 0.186 m/s for velocity, 1.573 degrees for pitch angle, and 0.139 meters for altitude when compared to the motion capture system. In comparison, the Crazyflie platform achieves RMSE values of 0.075 m/s, 1.619 degrees, and 0.021 meters for the same states during similar flight maneuvers. This shows that our system performs closely to the state-of-the-art Crazyflie in these scenarios. In future work, we aim to integrate TinySense with insect-scale robots to perform more advanced tasks, such as confined-space navigation [36].

### v2

@@
 Abstract
-In this paper, we investigate the prospects and challenges of sensor suites in achieving autonomous control for flying insect robots (FIRs) weighing less than a gram.
-FIRs, owing to their minuscule weight and size, offer unparalleled advantages in terms of material cost and scalability.
+In this paper, we introduce advances in the sensor suite of an autonomous flying insect robot (FIR) weighing less than a gram.
+FIRs, because of their small weight and size, offer unparalleled advantages in terms of material cost and scalability.
 However, their size introduces considerable control challenges, notably high-speed dynamics, restricted power, and limited payload capacity.
 While there have been notable advancements in developing lightweight sensors, often drawing inspiration from biological systems, no sub-gram aircraft has been able to attain sustained hover without relying on feedback from external sensing such as a motion capture system.
-The lightest vehicle capable of sustained hover—the first level of “sensor autonomy”—is the much larger 28 g Crazyflie.
+The lightest vehicle capable of sustained hovering—the first level of “sensor autonomy”—is the much larger 28 g Crazyflie.
 Previous work reported a reduction in size of that vehicle’s avionics suite to 187 mg and 21 mW.
@@
 A Kalman Filter (KF) fuses these measurements to estimate the state variables that are needed to control hover: pitch angle, translational velocity, and altitude.
-Our system achieved performance comparable to that of the Crazyflie’s estimator while in flight, with root mean squared errors of 1.573 degrees, 0.186 m/s, and 0.139 m, respectively, relative to motion capture.
+Our system achieved performance comparable to that of the Crazyflie’s estimator while in flight, with root mean squared errors of 1.573 deg, 0.186 m/s, and 0.136 m, respectively, relative to motion capture.
 I INTRODUCTION
@@
 Insect-scale robots have extreme size, speed, weight, and power (SSWaP) constraints.
-“Sensor autonomy” including stable hovering, represents first level of autonomy in the hierarchy proposed in [11], and is the foundation upon which higher level capability is built.
-We impose a requirement that sensing and computation to take no more than 10% of the power required to stay in the air; any more than this and flight time is severely impacted.
+“Sensor autonomy” including stable hovering, represents the first level of autonomy in the hierarchy proposed in [11], and is the foundation upon which higher level capability is built.
+We impose a requirement that sensing and computation take no more than 10% of the power required to stay in the air; any more than this and flight time is severely impacted.
 This factor holds for important examples of autonomous drones, such as the 1.5 kg system in [12] and the 30 g system in [13].
@@
 Figure 1: The TinySense sensor suite shown next to a U. Washington Robofly and a standard pencil for scale. The width of the sensor suite board is approximately 12 mm and its mass is 78.4 mg.
-The sensing system is very crucial for robots.
-Large drones can have more payloads to hold sensors onboard such as inertial measurement units (IMUs) and the global positioning system (GPS).
-However, GPS does not work well in indoor environments [14] and small robots are not able to hold many sensors due to constraints in size, speed, weight, and power consumption (SSWaP).
-There are several work done on the sensor suite for insect-scale robots  [15],  [6],  [16].
-The most recent sensor suite is 187 mg and consumes 21 mW [15].
-In this manuscript, we introduce a new avionics system for autonomous hovering flight that is even better tailored in mass and energy consumption for an insect-scale robot (Fig.
+Large drones can carry larger sensor payloads such as light-based ranging systems (LIDAR) and the global positioning system (GPS).
+However, GPS does not work well in indoor environments [14] and small robots are not able to hold emissive sensors due to constraints in size, speed, weight, and power consumption (SSWaP).
+Recent work has introduced a 187 mg avoinics suite that consumes 21 mW [15], considered gyroscope-free avionics [6], and has investigated the effect of wing-induced vibrations [16].
+In this paper, we introduce a new avionics system for autonomous hovering flight that is even better tailored in mass and energy consumption for an insect-scale robot (Fig.
 3) than previous work [15].
@@
 First, we replaced the power-hungry laser rangefinder with a pressure sensor.
-In recent years these have attained sufficient precision to be used for the purpose of altitude estimation and control even for insect-scale robots.
-Second, we replaced the optic flow sensor with a global shutter camera and compute optic flow on a microcontroller, resulting in a reduction in weight from 97 mg to 44 mg.
+In recent years, these have attained sufficient precision to be used for the purpose of altitude estimation and control, even for insect-scale robots.
+Second, we replaced the optic flow sensor with a global shutter camera and computed optic flow on a microcontroller, resulting in a reduction in weight from 97 mg to 44 mg.
 As in [15], this new sensor suite is able to estimate the crucial variables that are needed for controlled flight: pitch angle, translational velocity, and altitude.
@@
   1. 1.
-    A significant reduction in weight of the sensor suite to 78.4 mg which is less than half of the previous lightest sensor suite [15].
+    A reduction in weight of the sensor suite to 78.4 mg which is less than half of the previous lightest sensor suite [15].
   2. 2.
-    We showed that we were able to make accurate estimates of the state variables pitch angle, translational velocity, and altitude using data collected wirelessly from this sensor suite while it was mounted on a flying helicopter.
+    Demonstration of accurate estimates of the state variables pitch angle, translational velocity, and altitude using data collected wirelessly from this sensor suite while mounted on a drone.
   3. 3.
-    Implemented an optic flow algorithm on a microcontroller.
-Subsequent sections elucidate the dynamics, sensor configurations and measurement models, and corroborate our findings with data acquired wirelessly from our sensor suite, which is then analyzed on a desktop computer.
+    Optic flow estimation on a microcontroller.
+Subsequent sections elaborate on the dynamics, sensor configurations and measurement models, and validate our findings with data acquired wirelessly from our sensor suite, which is then analyzed on a desktop computer.
 II METHODOLOGY
@@
 Techniques to stabilize them passively through modification of wing configuration [20] or using air dampers [21] are either unproven or result in undesirable steady-state oscillations, respectively.
-To stabilize these dynamics, we will use a linearized dynamical model from which we will construct an observer of these unstable states so that they can be controlled.
+Our approach in this work is to first use linearized dynamics and observation models from which we can construct a Kalman observer of these unstable states so that they can be controlled.
 Figure 2: Forces and state variables of a small hovering Flying Insect Robot (FIR).
-Figure 2 depicts forces acting on a flapping-wing robot such as a gnat robot or the Robofly, but the proposed model can apply to almost any small flying aircraft, including EHD-based thrusters [22].
-Near hover, the vehicle’s dynamics are well approximated by a linear state-space model [23, 24, 21] of the form $\dot{\bm{q}}=A\bm{q}+B\bm{u}+G\bm{w}$, where $\bm{w}$ is the process noise, assumed to be roughly zero-mean Gaussian white noise with known covariance.
+Fig.
+2 depicts forces acting on a flapping-wing robot such as a gnat robot or the Robofly, but the proposed model can apply to almost any small flying aircraft, including EHD-based thrusters [22].
+Near hover, the vehicle’s dynamics are well approximated by a linear state-space model [23, 24, 21] of the form $\dot{\bm{q}}=A\bm{q}+B\bm{u}+G\bm{w}$, where $\bm{w}$ is the process noise, assumed to be zero-mean Gaussian white noise with known covariance.
 Process noise $\bm{w}$ is included because it is a necessary consideration for the design of the estimator.
 To simplify the analysis, we restrict our attention to motion in the $x$-$z$ plane; the dynamics in the $y$-$z$ plane are very similar.
-We choose a minimal state vector $\bm{q}=[\theta,v_{x},z]^{T}$, where $\theta$ is the pitch angle, $v_{x}$ is the vehicle’s translational velocity in the world $x$-direction, and $z$ is its vertical position or altitude (Fig.
-2).
-When the nonlinear Euler-Lagrange equation dynamics are linearized by taking the Jacobian at $\theta=0$, thrust force $f_{t}=mg$, and $z=z_{d}$, where $z_{d}$ is the desired altitude, the dynamics matrices are given by:
-$\footnotesize A=\left[\begin{matrix}0&0&0\\ g&-\frac{b}{m}&0\\ 0&0&0\end{matrix}\right],\quad B=\left[\begin{matrix}1\\ 0\\ 0\end{matrix}\right]$ (1)
-, where $m$ is the vehicle’s mass, and $b$ is the vehicles translational drag proportionality constant [16].
-Note that because we are concerned in this paper about only estimation, we ignore inputs from the motor, and instead use the measurements of the gyroscope $\bm{u}=[\omega_{m}]$ as the only input.
+We choose a minimal state vector $\bm{q}=[\theta,v_{x},z]^{T}$, where $\theta$ is the pitch angle (rad), $v_{x}$ is the vehicle’s translational velocity in the world $x$-direction (m/s), and $z$ is its vertical position or altitude (m) as illustrated in Fig.
+2.
+When the nonlinear Euler-Lagrange equation dynamics are linearized by taking the Jacobian at $\theta=0$, thrust force $f_{t}=mg$ (N), and $z=z_{d}$, where $z_{d}$ is the desired altitude (m), the dynamics matrices are given by:
+$A=\left[\begin{matrix}0&0&0\\ g&-\frac{b}{m}&0\\ 0&0&0\end{matrix}\right],\quad B=\left[\begin{matrix}1\\ 0\\ 0\end{matrix}\right],$ (1)
+where $m$ is the vehicle’s mass (kg), and $b$ is the vehicle’s translational drag proportionality constant (Ns/m) [24].
+Note that because we are concerned in this paper about only estimation, we ignore inputs from the motor, and instead use the measurements of the gyroscope $\bm{u}=[\omega_{m}]$ (rad/s) as the only input.
 This is discussed further in Section V.
@@
 The global positioning system is largely denied indoors and does not provide enough resolution anyway (1–10 m), and, further, requires significant power for signal processing.
-A lightweight and low-power sensing modality that can provide the necessary information to hover is “optic flow,” which is defined as the velocity of the image (in rad/s) across the camera imager [28].
+A lightweight and low-power sensing modality that can provide the necessary information to hover is “optic flow,” which is defined as the velocity of the image (in rad/s) across the camera imager [28, 29, 30].
 In tandem with other sensors, optic flow can be used to estimate the distance to obstacles as the imager translates through space [31].
 Methods for estimating optic flow include calculating luminance gradients locally [32] (Lucas-Kanade), or image-wide [33], as well as iterative search for matching blocks of pixels [34].
-Distances can also be estimated with a second camera and stereo vision using line matching [33], but this requires a second camera and a long “baseline” between cameras for accuracy, which would add too much weight.
-Figure 3: (a) Sensor suite (PAG7920LT optical sensor, ICM42688-P IMU, BMP388 pressure sensor) next to penny.
+Distances can also be estimated with a second camera and stereo vision using line matching [35], but this requires a second camera and a long “baseline” between cameras for accuracy, which adds excessive weight.
+Figure 3: Sensor suite (PAG7920LT optical sensor, ICM42688-P IMU, BMP388 pressure sensor) next to US 1 cent coin.
 Given these constraints, we propose the following set of onboard sensors: An optical sensor, a pressure sensor, and a gyroscope. We designed a customized microcircuit with the three sensors. (Fig. 3)
-III-A The camera and optic flow
-The system employs an optical sensor, the PAG7920LT (PixArt Imaging, Inc., Taiwan), to capture $160\times 120$ QQVGA images at 100 Hz over a 32 MHz SPI connection, with which the microcontroller computes optic flow via the Lucas-Kanade Method.
+III-A Camera and optic flow
+The system employs an optical sensor, the PAG7920LT (PixArt Imaging, Inc., Taiwan), to capture $160\times 120$ QQVGA images with a $120^{\circ}$ field of view at 100 Hz over a 32 MHz SPI connection, with which the microcontroller computes optic flow via the Lucas-Kanade Method.
 The reason we select this camera is that, to our knowledge, it is the only global shutter camera that provides an SPI interface to transmit the image data.
@@
 The Lucas-Kanade algorithm is performed by calculating the following quantities:
-$A=\left[\begin{matrix}I_{x}(q_{1})&I_{y}(q_{1})\\ I_{x}(q_{2})&I_{y}(q_{2})\\ ...&...\\ I_{x}(q_{n})&I_{y}(q_{n})\end{matrix}\right]\bm{v}=\left[\begin{matrix}V_{x}\\ V_{y}\end{matrix}\right]b=\left[\begin{matrix}-I_{t}(q_{1})\\ -I_{t}(q_{2})\\ ...\\ -I_{t}(q_{n})\end{matrix}\right]$
-, where $q_{1},q_{2},\dots,q_{n}$ are the brightness readings of each of the $n$ pixels in the image, $\bm{v}$ is the image flow in pixels per frame, and $I_{x}(q_{i}),I_{y}(q_{i}),I_{t}(q_{i})$ are the image’s horizontal, vertical, and temporal (time) partial derivatives respectively, at pixel $q_{i}$.
-Then, the optic flow is estimated (in pixels/frame) by taking the pseudoinverse $\bm{v}=(A^{T}A)^{-1}A^{T}b$.
+$J=\left[\begin{matrix}I_{x}(p_{1})&I_{y}(p_{1})\\ I_{x}(p_{2})&I_{y}(p_{2})\\ ...&...\\ I_{x}(p_{n})&I_{y}(p_{n})\end{matrix}\right],\bm{v}=\left[\begin{matrix}V_{x}\\ V_{y}\end{matrix}\right],b=\left[\begin{matrix}-I_{t}(p_{1})\\ -I_{t}(p_{2})\\ ...\\ -I_{t}(p_{n})\end{matrix}\right],$
+where $p_{1},p_{2},\dots,p_{n}$ are the pixels in the image, $\bm{v}$ is the image flow in pixels per frame, and $I_{x}(p_{i}),I_{y}(p_{i}),I_{t}(p_{i})$ are the image’s horizontal, vertical, and temporal (time) partial derivatives respectively, at pixel $p_{i}$.
+Then, the optic flow is estimated (in pixels/frame) by taking the pseudoinverse $\bm{v}=(J^{T}J)^{-1}J^{T}b$.
 To minimize computation for the limited microcontroller resources, we reduced the image resolution by a factor of 4 (producing a $40\times 30$ resolution) by skipping pixels.
-The two most expensive operations involve computing $A^{T}A$ and $A^{T}b$.
-A $40\times 30$ image produces an $A$ matrix with $(40-2)*(30-2)=1064$ rows and 2 columns and a $b$ matrix with 1064 rows and 1 column.
-We estimate multiplying two matrices of size $k\times l$ and $l\times m$ will require $kml$ multiplications.
-Thus, computing $A^{T}A$ and $A^{T}b$ require 4256 and 2128 multiplications, respectively.
+The two most expensive operations involve computing $J^{T}J$ and $J^{T}b$.
+A $40\times 30$ image produces an $J$ matrix with $(40-2)\times(30-2)=1064$ rows and 2 columns and a $b$ matrix with 1064 rows and 1 column.
+Multiplying two matrices of size $k\times l$ and $l\times m$ requires at worst $kml$ multiplications.
+Thus, computing $J^{T}J$ and $J^{T}b$ require 4256 and 2128 multiplications, respectively.
 We neglect the additional operations required to compute their product (2 multiplications) and the $2\times 2$ matrix inverse (6 multiplications and one divide), giving a total operations count of 6384.
 We additionally implemented a form of sparse optic flow in which flow vectors were computed for twelve $10\times 10$ patches, which are averaged to produce a final estimate.
-A $10\times 10$ image produces an $A$ matrix with $(10-2)*(10-2)=64$ rows and 2 columns and a $b$ matrix with 64 rows and 1 column.
-Thus, computing $A^{T}A$ and $A^{T}b$ require approximately 256 and 128 multiplications, respectively.
-With twelve matrices of size $10\times 10$, the total number of multiplication is $12(256+128)=4608$.
-This is a reduction of approximately $39\%$.
-In the future, we will evaluate these patches individually to remove outliers, removing patches with low eigenvalues (low contrast) or significantly different eigenvalues (aperture problem).
-To convert the image flow in pixels/frame to angular velocity in radians/s, we rotated the TinySense by hand and minimized the error between the optic flow measurement multiplied by a constant $\alpha$ and the gyroscope reading.
-This constant, which depends on the pixel density and field of view of the camera, is given by $\alpha=-0.027$ rad/pixel.
+A $10\times 10$ image produces an $J$ matrix with $(10-2)\times(10-2)=64$ rows and 2 columns and a $b$ matrix with 64 rows and 1 column.
+Thus, computing $J^{T}J$ and $J^{T}b$ require approximately 256 and 128 multiplications, respectively.
+With twelve matrices of size $10\times 10$, the total number of multiplication is $12\times(256+128)=4608$.
+This is a reduction of approximately $28\%$.
+To convert the optic flow from pixels/frame to angular velocity (rad/s), we found the calibration constant $\alpha=-0.027$ (rad/pixel) by rotating the TinySense by hand and performing a linear regression against its gyroscope readings.
 III-B Pressure sensor
-Altitude measurements are conducted by a Bosch BMP388 pressure sensor.
-This sensor estimates altitude as a proportion of the difference between the surrounding air pressure and sea level.
+Altitude measurements were conducted by a Bosch BMP388 pressure sensor.
+This sensor estimates altitude as a proportion of the difference between the air pressure and sea level.
 We set it for 2$\times$ pressure oversampling and an IIR filter coefficient of 3 to maximize resolution.
 To compensate for variations in local pressure causing inaccurate altitude measurements, the sensor was assigned a bias equal to the average of the first 25 measurements where the sensor system was grounded; this bias was subtracted from every reading.
-III-C The gyroscope
+III-C Gyroscope
 For the onboard inertial measurement unit (IMU), we chose the TDK ICM42688-P because it is a recent model that is available in the smallest commercially-available package size of 2.5 $\times$ 3 mm.
@@
 Note that the gyroscope provides a low-noise measurement of angular velocity $\bm{\omega}$.
-In cases like this, where the sensor provides more precise information than can be derived from knowing the system’s actuated inputs (e.g.
-the torque applied by the aircraft’s rotors or wings), it is common practice to re-formulate the Kalman Filter to use the sensor’s readings as its "input" $\bm{u}$, rather than the actual system input.
-Doing this has the additional advantage that for each such measurement, one less state variable is required, reducing computation requirements and complexity.
-In this case, it means that the angular velocity $\omega$ need not appear among the state variables.
-The following is our observation model for the suite of sensors, which is:
+In cases like this in which the sensor provides more precise information than can be derived knowing the system’s actuated inputs (e.g.
+torque applied by the aircraft’s rotors or wings), it is common practice to reformulate the Kalman Filter to use the sensor’s readings as “input” $\bm{u}$, and to ignore actual system input.
+This has an additional advantage that one less state variable is required for each such measurement.
+Our system includes gyroscope readings as one of its inputs and hence angular velocity $\omega$ does not appear among the state variables.
+Our observation model for the suite of sensors is:
 $\bm{y}=\left[\begin{matrix}\Omega_{m}+n_{o}\\ z_{m}+n_{p}\\ \end{matrix}\right],$ (2)
-where $\Omega_{m}$ is the optic flow measurements from optic flow sensor, $z_{m}$ is the measurement from the pressure sensor, and $n_{o}$ and $n_{p}$ are the noise terms for the two sensors.
+where $\Omega_{m}$ is the optic flow measurements (rad/s) from optic flow sensor, $z_{m}$ is the measurement (m) from the pressure sensor, and $n_{o}$ and $n_{p}$ are the noise terms for the two sensors.
 We assume noise terms are zero-mean Gaussian white noise, a standard assumption in sensor noise modeling to simplify the analysis.
-The measurement model for the optic flow in the $x$ direction, as measured by the optic flow camera is a nonlinear function of state [6]:
+The measurement model for the optic flow in the $x$ direction, as measured by the optic flow camera is a nonlinear function of the state [6]:
 $\Omega_{m}=\omega_{m}-\frac{v_{x}}{z},$ (3)
-where $\omega_{m}$ is the angular velocity measurement from gyroscope. We can see the optic flow measurement depends on $z$. We can take the Jacobian linearization at a desired altitude $z_{d}$:
+where $\omega_{m}$ is the angular velocity measurement (rad/s) from gyroscope. The optic flow measurement depends on $z$. We can take the Jacobian linearization at a desired altitude $z_{d}$:
 $\begin{split}\Omega_{m}(z)=\Omega_{m}(z_{d})+O(z)=\omega_{m}-\frac{v_{x}}{z_{d}}+O(z).\end{split}$ (4)
-With this linearization, our measurement model can be cast into the traditional Kalman Filter model, $\bm{y}=C\bm{q}+D\bm{u}+\bm{n}$, where
+With this linearization, our measurement model can be cast into the Kalman Filter model, $\bm{y}=C\bm{q}+D\bm{u}+\bm{n}$, where
 $C=\left[\begin{matrix}0&-\frac{1}{z_{d}}&0\\ 0&0&1\end{matrix}\right],\quad D=\left[\begin{matrix}1\\ 0\end{matrix}\right].$ (5)
 The observability matrix [$C$; $CA$] is full rank, meaning that a Kalman Filter will be able to estimate the system’s state using the measurements given in Eq. (5).
-The Kalman Filter observer estimates $\hat{\bm{q}}$ by numerically integrating the system
+The Kalman Filter observer estimates $\hat{\bm{q}}$ by numerically integrating the system:
 $\dot{\hat{\bm{q}}}=A\hat{\bm{q}}+B\bm{u}+K(\bm{y}-C\hat{\bm{q}}-D\bm{u}),$ (6)
 where $K$ is the Kalman gain.
-We used a gradient descent algorithm to minimize the RMSE between the TinySense state estimates and the mocap ground-truth to tune the process noise covariance matrix $Qn$ and sensor noise covariance matrix $Rn$:
-$Qn=diag(0.14^{2},0.093^{2},0.08^{2}),\quad Rn=diag(1.06^{2},0.2^{2}).$
+We calculated optic flow variance $R_{n1}$ to be 0.017 by subtracting its value from that predicted by mocap using Eq.
+(3) during a constant-velocity horizontal translational flight.
+We calculated the variance of the pressure sensor $R_{n2}$ to be 0.0055 by taking measurements while holding it at a constant altitude for 60 seconds:
+$R_{n}=\mathrm{diag}(R_{n1},R_{n2})=\mathrm{diag}(0.017,0.0055).$
+To estimate the process noise covariance matrix $Q_{n}$, we employed a hybrid optimization approach, combining grid search for coarse parameter initialization, gradient descent for fine-tuning, and manual adjustment to minimize the RMSE between TinySense state estimates and mocap ground truth:
+$Q_{n}=\mathrm{diag}(0.0124^{2},0.001^{2},0.22^{2}).$
 For the process noise (disturbance) matrix $G$, we assumed that the noise enters the system as white noise disturbance adding to the derivatives of the three states $\theta$, $v_{x}$, and $z$.
-Based on these, the Kalman gain $K$ was computed using the lqe command in python-control [34].
-$G=\left[\begin{matrix}0.72&0&0\\ 0&0.10&0\\ 0&0&20.00\end{matrix}\right],\quad K=\left[\begin{matrix}0.00204&0\\ 0.48452&0\\ 0&1.81818\end{matrix}\right].$
+Based on these, the Kalman gain $K$ was computed using the lqe command in python-control [36]:
+$G=\left[\begin{matrix}1&0&0\\ 0&1&0\\ 0&0&1\end{matrix}\right],\quad K=\left[\begin{matrix}0.095&0\\ -1.32&0\\ 0&3\end{matrix}\right].$
 We implemented the Kalman Filter offline in Python by iteratively computing the state estimates $\hat{\bm{q}}_{i+1}$ according to $\hat{\bm{q}}_{i+1}=\hat{\bm{q}}_{i}+\dot{\hat{\bm{q}}}_{i}\cdot dt$.
-Figure 4: Sensor measurements from the onboard gyroscope, optic flow sensor, and pressure sensor of TinySense are compared with those from the Crazyflie for three different flight experiments (a), (b), and (c).
+Figure 4: Sensor measurements from the onboard gyroscope, optic flow sensor, and pressure sensor of TinySense are compared with those from the Crazyflie for three different flight experiments (a), (b), and (c); trajectories are described in the main text and corresponding state estimates in Fig.
+5.
 In the altitude plots, altitude is measured by the Crazyflie’s laser rangefinder, while the TinySense uses a pressure sensor; mocap data is also included for comparison.
@@
 V-B Motion capture system
-To provide a ground-truth comparison for the sensors, we used a twelve-camera OptiTrack Flex 13 motion capture arena.
+To provide a ground-truth comparison for the sensors, we used a twelve-camera OptiTrack Flex 13 motion capture arena, with spatial accuracy $\pm$ 0.20 mm.
 The Crazyflie drone was fitted with four retro-reflective motion capture markers and its position and orientation were recorded at 120 Hz.
@@
 4.
-Figure 5: State estimates from TinySense after Kalman filtering, Crazyflie estimates, and mocap estimates are compared.
+Figure 5: Comparison between state estimates from TinySense’s Kalman Filter, Crazyflie, and mocap.
 We conducted three flight experiments (a), (b), and (c), with the Kalman filter estimates starting at time zero.
-Note that we ignore the pressure sensor measurements in the first 1.8s because of the pressure change near the surface.
-TABLE II: Estimation of Optic Flow Algorithm Power Use
-Update Frequency (Hz) Occurences/ Update Step Name Single-cycle Operations Int Division Operations Floating-point Division Operations Total Cycles/Update Total Cycles/Second Power Use
-100 1 Skip frame 8493 0 8 249781 24.9781 Mhz $\sim$5.4552mW
+Pressure sensor measurements were ignored during the first 1.8 s because pressure changes due to the ground effect at takeoff confound readings.
+TABLE II: Estimate of optic flow algorithm power use
+update frequency (Hz) occurences/ update task single-cycle operations int division operations floating-point division operations total cycles/update total cycles/s (Mhz) estimated power (mW)
+100 1 skip frame 8493 0 8 249781 25.0 5.5
 12 LK OF 18434 128 2
@@
 V-D Computational load
-We estimated the power consumption of the optic flow algorithm from the number of cycles in a single run.
-We focus on only estimating three types of operations conducted on the ARM Cortex-M4 based microcontroller: single-cycle operations (addition, subtraction, multiplication, and memory read/write); integer divisions, which require 12 cycles at most; and floating-point divisions, which require 14 cycles.
-These results are summarized in Table II.
-Given that in our current system, all sensors update at a rate of 100 Hz, we gather that the total number of cycles per second is $\sim$24.98 Mhz.
-Operating on the nRF52840 chip with an efficiency of 52 $\mu$A/Mhz and a battery-powered voltage of 4.2V, we estimate the power use of the optic flow algorithm to be $\sim$5.45mW.
-V-E Results and Discussions
+Table II provides an estimate of the power consumption of the optic flow algorithm, by calculating the number of clock cycles needed for each operation.
+Our estimate includes the matrix operations described in Section III-A, as well as additional housekeeping operations, and is based on ARM Cortex-M4 instruction timings for addition, subtraction, multiplication, and memory read/write operations; integer divisions require 12 cycles at most and floating-point divisions require 14 cycles.
+Given that in our current system all sensors update at a rate of 100 Hz, we estimated that the total number of needed cycles per second is $\sim$25.0 Mhz.
+Operating the nRF52840 with an efficiency of 52 $\mu$A/MHz and a battery-powered voltage of 4.2 V, we estimate the power usage of the optic flow algorithm to be 5.5 mW.
+V-E Results and discussion
 Fig.
-5 compares our sensor suite’s state estimate to that of the Crazyflie’s built-in estimator, and states derived from the mocap system, considered to be ground truth.
+5 compares our sensor suite’s state estimate to that of the Crazyflie’s built-in estimator.
+States derived from the mocap system are considered to be ground truth.
 For the state estimator, time $t=0$ corresponds to when the gyroscope reading becomes significantly different from zero, indicating liftoff.
-It can be observed in Fig.
-4 that a pressure increase occurs when the propellers first turn on, which can be observed as a drop in measured altitude.
+Fig.
+4 shows that a pressure increase occurs when the propellers first turn on, which can be observed as a drop in measured altitude.
 To mitigate unwanted effects of this in the Kalman Filter, we ignored (set to zero) altitude measurements $z_{m}$ during the first 1.8 s after liftoff.
 Nevertheless, once it is in the air, the altitude measurement matches quite closely with the laser rangefinder on the Crazyflie, though with slightly more noise.
-Our system’s estimate compares well with those of the Crazyflie’s, despite receiving gyroscope readings at a much lower rate (100 Hz) compared to 1000 Hz on the Crazyflie[35].
-The RMSE is approximately equal for pitch angle (Table III).
-We observe that there are occasionally large errors in translational velocity $v_{x}$.
-These appear to occur after sudden changes in attitude that occur around $t=5$ s and $t=6$ s, during which the angular velocity momentarily exceeds 3 rad/s (Fig.
+Our system’s estimate compares well with the Crazyflie’s, despite receiving gyroscope readings at only 100 Hz, which is ten times slower [37].
+The key metric of pitch angle RMSE is approximately equal (Table III).
+Large errors in translational velocity $v_{x}$ appear around $t=5$ and $t=6$.
+These occur after sudden changes in attitude, during which angular velocity exceeds 3 rad/s (Fig.
 4).
-We note that at these moments, our TinySense optic flow sensor’s readings are consistently lower than those of the Crazyflie.
-Lucas-Kanade cannot detect motion accurately if the displacement between frames is greater than a half pixel between image frames.
-The maximum allowable angular velocity is given by $\Omega=0.5\alpha F=(0.5)(0.027)(100)=1.4$ rad/s, where $\alpha$ is the angular increment between neighboring pixels (Section II-A) and $F$ is the frame rate in Hz.
-So the expected optic flow is faster than can be detected.
-This could be mitigated by a faster frame rate or a larger angular displacement between pixels.
+At these moments, we observed that TinySense optic flow readings are consistently lower than those of the Crazyflie.
+Lucas-Kanade cannot detect motion accurately if the displacement between frames is greater than a half pixel between image frames [38].
+This limits optic flow measurement to $\Omega_{m}<0.5\alpha F=1.4$ rad/s, where $\alpha$=0.027 is the angular increment between pixels and $F$=100 is the frame rate (Hz).
+We believe that the lateral velocity estimate diverges during these fast maneuvers because it depends on erroneous optic flow readings.
 TABLE III: Root Mean Squared Error (RMSE) of estimates relative to mocap system
@@
 Crazyflie 0.075 $\pm$ 0.009 1.619 $\pm$ 0.267 0.021 $\pm$ 0.001
-TinySense 0.186 $\pm$ 0.015 1.573 $\pm$ 0.166 0.139 $\pm$ 0.026
+TinySense 0.186 $\pm$ 0.015 1.573 $\pm$ 0.166 0.136 $\pm$ 0.026
 VI Conclusion
 In the present study, we introduce a new sensor suite with dramatically reduced mass, and associated optic flow and state estimation software that is compatible with the computational constraints of an onboard microcontroller small enough to fly aboard flying insect robots (FIRs) weighing less than a gram.
-Our proposed avionics suite diverges in two important ways compared to that of the 28 g Crazyflie helicopter, the lightest vehicle yet to perform controlled hovering, and previous work [15].
-First, it replaces a power-hungry laser rangefinder with a much more efficient (and slightly less ideal) and lighter pressure sensor.
-Second, we replaced an off-the shelf optic flow sensor, which is available only in relatively heavy packages, with a custom camera with a global shutter and a custom-written Lucas-Kanade-based optic flow algorithm running onboard a 10 mg microcontroller small enough to fly onboard an FIR (estimated conpute power usage is about 5 mW).
-Our system uses a Kalman Filter to estimate pitch angle, translational velocity, and altitude.
-Remarkably, the cumulative weight of this sensor suite is only 78.4 mg, rather less than the 187 mg of previous work [15].
-Incorporating a microcontroller to this configuration, the overall system remains well within the anticipated 252 mg payload limit of the 143 mg robotic platform reported in [18].
-Our system demonstrates average RMSE values of 0.186 m/s for velocity, 1.573 degrees for pitch angle, and 0.139 meters for altitude when compared to the motion capture system.
-In comparison, the Crazyflie platform achieves RMSE values of 0.075 m/s, 1.619 degrees, and 0.021 meters for the same states during similar flight maneuvers.
+Our proposed avionics suite diverges in two important ways from the 28 g Crazyflie helicopter, the lightest vehicle yet to perform controlled hovering, and previous work [15].
+First, we replaced a power-hungry laser rangefinder with a much more efficient (and slightly less ideal) and lighter pressure sensor.
+Second, we replaced an off-the-shelf optic flow sensor, which is available only in relatively heavy packages, with a custom camera with a global shutter and a custom-written Lucas-Kanade-based optic flow algorithm running onboard a 10 mg microcontroller small enough to fly onboard an FIR.
+We estimate the compute power usage is about 5 mW.
+Our system uses a linear Kalman Filter to estimate pitch angle, translational velocity, and altitude.
+This can be readily adapted to a nonlinear extended Kalman Filter for more generality [15].
+The cumulative weight of this sensor suite is only 78.4 mg, rather less than the 187 mg of previous work [15].
+With the addition of a 30 mg microcontroller and crystal, the overall system remains well within the 252 mg payload limit of the 143 mg robotic platform reported in [20].
+Our system demonstrates average RMSE values of 0.186 m/s for velocity, 1.573 deg for pitch angle, and 0.136 m for altitude when compared to the motion capture system.
+In comparison, the Crazyflie platform achieves RMSE values of 0.075 m/s, 1.619 deg, and 0.021 m for the same states during similar flight maneuvers.
 This shows that our system performs closely to the state-of-the-art Crazyflie in these scenarios.
-In future work, we aim to integrate TinySense with insect-scale robots to perform more advanced tasks, such as confined-space navigation [36].
+The pressure sensor showed negligible drift over a 60-second period, providing a usable altitude estimate as long as its output during takeoff is ignored.
+Future improvements include mitigating the lateral velocity errors that occur during rapid rotations that exceed the limit of Lucas-Kanade estimation (Section V-E) by increasing the frame rate or image blurring.
+Additionally, our patch-based optic flow estimation (Section III-A) may allow outliers to be rejected based on their eigenvalues to eliminate low-contrast patches, reducing noise [38].
+Future work will integrate TinySense along with a microcontroller and power hardware onto insect-scale robots such as the Robofly [39] to perform visual navigation [40].
+The Robofly performs faster accelerations than the Crazyflie and is a different vibration environment, but is otherwise similar.
+Previous work shows vibrations do not confound gyroscope readings [41, 16].
+Algorithmic changes in optic flow may be needed to accommodate vibration, but our choice of a global shutter camera will mitigate tearing artifacts.